# Latent Inversion with Timestep-aware Sampling for Training-free Non-rigid Editing

### v2

@@ 4 Method
We follow the text optimization from Imagic to preserve the original identity and achieve smooth non-rigid editing (See Figure 3(A)). As text optimization alone is insufficient for a perfect match with the input image, Imagic solves this problem via model finetuning. However, in the Stable Diffusion setting, we found that the model finetuning presents several significant drawbacks that adversely affect editing quality. Firstly, it triggers a severe tradeoff between (1) overfitting, where the model overly memorizes the original image and thus cannot faithfully incorporate the text prompt well, or (2) the lack of identity preservation, where the edited image aligns with the target text but it fails to preserve the original image’s identity. Figure 2 provides visual evidence that the edited image closely resembles the input image as we finetune the model longer. Conversely, minimal finetuning produces an arbitrary image that bears little resemblance to the input image, indicating the lack of identity preservation (Text optim step 200 + Finetune step 100, +Finetune step 400 in Figure 2). Model finetuning thus hardly satisfies both identity preservation and editability simultaneously. (More ablation experiments on finetuning steps in the Appendix.) Even worse, it causes color distortion, disrupting the data distribution. This is because model training aims to reach the data distribution of the training dataset, which in this context consists of only a single input image.
@@
While employing the inverted noise helps maintain the overall color and context of the original image, introducing a new text prompt different from the source text can cause deviations in image configuration, such as the size and position of the object, compared to the original image. To address this issue, we leverage a well-known observation that diffusion models generate global structures at the initial timesteps and fine details toward the end Choi et al. (2022). Specifically, we devise an additional regularization depending on sampling timesteps (See Figure 3(C)). During the initial steps ($n\rightarrow T$), we incorporate the source text embedding and then transition to interpolated embeddings for the subsequent timesteps:
@@
This strategy helps to closely adhere to the global structure of the original image, a key component of the image identity, during the early stage of the editing process. Since then, the editing concept has been applied to the subsequent sampling steps, effectively incorporating editing concepts without altering the image identity. We show its substantial impact on preserving the identity through ablation studies in Section 5.4.
Figure 4: Qualitative results. We compare our editing results to five recent editing methods. While other methods suffer from either overfitting the input image or lacking identity preservation, our method shows the best structure preservation and edit fidelity.
@@ 5.1 Implementation Details
Dataset. We employ the Textual Editing Benchmark (TEdBench), introduced by Imagic Kawar et al. (2023). It comprises input images with target texts describing desired complex non-rigid editing.
@@
Baselines. For a comprehensive comparison, we evaluate our method against the following competitive techniques: Null-text inversion Mokady et al. (2023), Edit Friendly Huberman-Spiegelglas et al. (2023), Imagic Kawar et al. (2023), DDS Hertz et al. (2023), and Masactrl Cao et al. (2023). Notably, Imagic and Masactrl are specialized for the non-rigid editing task. Null-text inversion and Edit-Friendly represent prominent methods designed for natural image editing via inversion. DDS proposes a new paradigm of image editing using score distillation. In our evaluations, we run the released codes of all baseline models using pre-trained Stable Diffusion v1.4 and anime-style model Anything-V4 with publicly available checkpoints. Unfortunately, we could not include other non-rigid editing methods, such as LayerDiffusion Li et al. (2023) and KVInversion Huang et al. (2023a), as their codes are not released as of our submission date. For sampling, we use a conventional setting of DDIM with 50 timesteps  Cao et al. (2023); Huang et al. (2023a); Li et al. (2023) for all methods. For our method using DDIM inversion Song et al. (2020), we used a guidance scale of 1 for both inversion and sampling stages, as it is known that using a larger guidance scale leads to suboptimal results. For our method using Null-text inversion Mokady et al. (2023), we used a guidance scale of 7.5 for both inversion and sampling following previous works Cao et al. (2023); Huang et al. (2023a).
Evaluation Metrics. For measuring target text alignment and input image identity preservation quality of edited images, we utilize CLIP-T score Radford et al. (2021) and LPIPS score Zhang et al. (2018), following conventions Kawar et al. (2023); Mokady et al. (2023). CLIP score concerning the target text measures the cosine similarity between the embeddings of the target text prompt and the edited image. LPIPS score measures the difference between the original input image and the edited image in the VGG Simonyan and Zisserman (2014) feature space. However, these two metrics are incapable of evaluating the aesthetic quality of the edited image; they cannot penalize the distorted or unnatural outputs. To mimic the standard of human visual inspection, we additionally employ an Aesthetic score Schuhmann (2022), which measures how much people like the image on average. Few MLP layers are added on the CLIP ViT-L/14, and only the MLP layers are finetuned to predict the aesthetic score using MSE or MAE loss. AVA Murray et al. (2012) and Simulacra Pressman et al. (2022) are used for training the predictor, which are datasets that collect human ratings of 1 to 10 on real and generated images each. This metric is used to clean the LAION-5B Schuhmann et al. (2022) dataset.
5.2 Comparison with Other Works
@@
Results with Stable Diffusion. We compare the editing results of our method to other baselines using a pre-trained Stable Diffusion v1.4 in Figure 4. We show our results using Null-text inversion as it provides the best editing performance (Figure 7). For a fair comparison, we also applied Null-text inversion in Masactrl, which originally used DDIM inversion. It results in improved identity preservation and editability of Masactrl. We tried various hyperparameters of each method and chose the best result for each.
@@
Firstly, Null-text inversion alone for editing, in which we just input an inverted image and a target text through the model, shows low identity preservation. It does not retain the original structure of the input image and simply pursues generating an image that aligns well with the text prompt. This is expected because there is no other constraint for identity preservation during the sampling process. Meanwhile, Imagic has a serious tradeoff between overfitting and editability. Model finetuning especially causes a severe overfitting problem, which distorts the data distribution, whereas a random noise without finetuning during the sampling process ruins the structure preservation. Despite the extensive hyperparameter search, it is hard to find a sweet spot between these two factors (See Appendix). Besides, DDS also suffers from overfitting to the input image, especially when handling non-rigid editing. This shows that score distillation has trouble changing geometry. On the other hand, Edit-Friendly and Masactrl attempt to apply an editing concept while maintaining the original identity by improving latent noise space and utilizing attention maps, respectively. When an image has no complicated background and contains a single object, such as a Spiderman image in the third row of Figure 4, their editing quality is sufficiently high. However, they show unnatural edited results in other cases. In the first row of Figure 4, a bulldog opens its mouth but does not smile in Edit-Friendly, and the bulldog shows a grotesque mouth in Masactrl. In the second row, a horse jumps in both methods, but the hind legs of the horse remain similar to the input image, resulting in an unnatural jumping appearance. Overall, our method clearly shows the editing concept while faithfully preserving the original identity, like a bulldog with a big smile and a horse jumping dynamically. These visual trends and the superiority of our method consistently hold in other results. More qualitative results are demonstrated in Appendix.
Results with Anything-V4. We also perform our method on a pretrained anime-style model, Anything-V4. We generated an image using the model and then performed editing just like the real image editing setting. Figure 5 shows various editing results using diverse text prompts. We can change the posture, facial expression, and viewpoint smoothly while maintaining the original structure of the input image. This proves that our method generalizes well regardless of the image domain. More qualitative results are shown in Appendix.
Figure 5: Anything-v4. This figure showcases the adaptability of our model on an Anything-V4 anime-style model. Our method effectively edits the input image’s posture and viewpoint with various text prompts, smoothly preserving the original structure.
@@ 5.3 Quantitative Evaluation
We conduct a quantitative evaluation using the following non-rigid editing models: Imagic, Masactrl, and the proposed method. The experiment is carried out using the pre-trained Stable Diffusion v1.4. To ensure a fair comparison, we use Null-text inversion for both ours and Masactrl. We utilized the TedBench dataset to generate 86 edited images, producing five variants for each edited image with different seeds, resulting in a total of 430 images. In Figure 6, we present CLIP and Aesthetic scores of each method at different LPIPS scores, where LPIPS changes depending on hyperparameters. Herein, a higher CLIP score refers to better alignment with target text while keeping the LPIPS score fixed. Imagic consistently yields the lowest CLIP score across all settings, indicating that its editability stays far behind ours and Masactrl. Our method exhibits the highest CLIP score in most settings. It implies that our method effectively preserves the input image’s identity while accurately conveying the editing concept compared to other methods. Regarding the Aesthetic score, our method achieves the highest score in all the settings. It highlights that our method has both superior editability and aesthetic appeal compared to other methods. This success can be attributed to our inversion and timestep-aware text injection sampling, which preserve the original identity without disrupting the data distribution. This enables text optimization to smoothly transition towards the desired editing concept while maintaining the background and object identity faithfully. Imagic tends to overfit the input image, maintaining the naturalness of the image structure, whereas color gets distorted. Consequently, its aesthetic score lags behind ours, but not the least. Finally, Masactrl presents the lowest Aesthetic score, consistent with its tendency to generate unnatural images with distorted object appearances and poses. This can be attributed to the difference in object appearance and poses between the source and target images, each generated using the original source and target text, respectively. Example images are available in the Appendix. As the appearance gap between the target image and the source image intensifies, inserting the attention maps of the source image into the target image sampling process often results in abnormal objects with unnatural appearances and poses. Though there is a point where Masactrl has a higher CLIP score compared to ours, its low Aesthetic score shows that it has difficulty generating an edited image in which the editing concept seamlessly blends with the object’s intrinsic characteristics.
Figure 6: Quantitative results. The graph represents (A) CLIP scores ($\uparrow$), indicating alignment with target text, and (B) Aesthetic scores ($\uparrow$), under various image quality levels as LPIPS scores change. Our method outperforms Imagic and Masactrl in both CLIP and Aesthetic scores in most settings, demonstrating superior editability and aesthetic quality. This highlights the effectiveness of our method, latent inversion with timestep-aware text injection sampling, in preserving the original identity and data distribution, leading to more natural and concept-aligned edits.
@@ 5.4 Ablation Experiments
We show the effectiveness of each component of our method through ablation experiments on inversion type and source text injection ratio.
Inversion Type. Figure 7 demonstrates the impact of different inversion methods: DDIM Song et al. (2020), Null-text inversion Mokady et al. (2023) with (NTI-src) and without source prompt (NTI-w/o src). Each image is edited with the target text “A photo of a jumping dog.” When employing DDIM inversion for our method, it is highly close to the input image but has low editability. This again proves that DDIM inversion is suboptimal for real image editing applications. On the other hand, Null-text inversion is known to improve inversion and editing quality, particularly when using classifier-free guidance with a guidance scale greater than 1, as compared to DDIM inversion. Interestingly, when using the source prompt, the preservation quality sometimes significantly deteriorates while focusing on editing (NTI - src). Conversely, when we set a source text prompt equal to null text, the edited images show significant improvement in structural preservation (NTI - w/o src). Therefore, when the reconstruction performance of inversion falls behind, employing the null text for the source prompt can enhance the overall reconstruction ability. Overall, Null-text inversion demonstrates superior performance in editing quality to DDIM.
Figure 7: Ablation on inversion type. We demonstrate our editing results with three inversion types: DDIM, Null-text inversion with source (NTI-src), and without source embedding (NTI-w/o src). The results confirm that DDIM is less favorable to the editing task than NTI, where setting null text instead of source text often produces more plausible results.

### v3

@@ 4 Method
We follow the text optimization from Imagic to preserve the original identity and achieve smooth non-rigid editing (See Figure 2 (A)). As text optimization alone is insufficient for a perfect match with the input image, Imagic solves this problem via model finetuning. However, in the Stable Diffusion setting, we found that the model finetuning presents several significant drawbacks that adversely affect editing quality. Firstly, it triggers a severe tradeoff between (1) overfitting, where the model overly memorizes the original image and thus cannot faithfully incorporate the text prompt well, or (2) the lack of identity preservation, where the edited image aligns with the target text but it fails to preserve the original image’s identity. Figure 1 provides visual evidence that the edited image closely resembles the input image as we finetune the model longer. Conversely, minimal finetuning produces an arbitrary image that bears little resemblance to the input image, indicating the lack of identity preservation (Text optim step 200 + Finetune step 100, +Finetune step 400 in Figure 1). Model finetuning thus hardly satisfies both identity preservation and editability simultaneously. (More ablation experiments on finetuning steps are in Appendix.) Even worse, it causes color distortion, disrupting the data distribution. This is because model training aims to reach the data distribution of the training dataset, which in this context consists of only a single input image.
@@
While employing the inverted noise during sampling process helps maintain the overall color and context of the original image, introducing a new text prompt different from the source text can cause deviations in image configuration, such as the size and position of the object, compared to the original image. To address this issue, we leverage a well-known observation that diffusion models generate global structures at the initial timesteps and fine details toward the end [Choi et al.(2022)Choi, Lee, Shin, Kim, Kim, and Yoon]. Specifically, we introduce an additional regularization depending on sampling timesteps (See Figure 2 (C)). During the initial steps ($n\rightarrow T$), we incorporate the source text embedding and then transition to interpolated embeddings for the subsequent timesteps:
@@
This strategy helps to closely adhere to the global structure of the original image, a key component of the image identity, during the early stage of the editing process. Since then, the editing concept is applied to the subsequent sampling steps, effectively incorporating editing concepts without altering the image identity. We show its substantial impact on preserving the identity through ablation studies in Section 5.4.
@@ 5.1 Implementation Details
Baselines. For a comprehensive comparison, we evaluate our method against the following competitive techniques: Imagic [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani], Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or], DDS [Hertz et al.(2023)Hertz, Aberman, and Cohen-Or], Masactrl [Cao et al.(2023)Cao, Wang, Qi, Shan, Qie, and Zheng] and Edit-Friendly [Huberman-Spiegelglas et al.(2023)Huberman-Spiegelglas, Kulikov, and Michaeli]. Notably, Imagic and Masactrl are specialized for the non-rigid editing task. Null-text inversion and Edit-Friendly represent prominent methods designed for natural image editing via inversion. DDS proposes a new paradigm of image editing using score distillation. In our evaluations, we run the released codes of all baseline models using pre-trained Stable Diffusion v1.4 with publicly available checkpoints. Unfortunately, we could not include other non-rigid editing methods, such as LayerDiffusion [Li et al.(2023)Li, Huang, Ding, and Li] and BARET [Qiao et al.(2024)Qiao, Wang, Su, Zhang, Yu, Wu, and Qi], as their codes are not released as of our submission date. Details on experiments are in Appendix.
Evaluation Metrics. We utilize CLIP-T score [Radford et al.(2021)Radford, Kim, Hallacy, Ramesh, Goh, Agarwal, Sastry, Askell, Mishkin, Clark, et al.] for measuring target text alignment and LPIPS [Zhang et al.(2018)Zhang, Isola, Efros, Shechtman, and Wang], PSNR, and MSE for measuring input image identity preservation quality of edited images. However, these metrics are incapable of evaluating the aesthetic quality of the edited image; they cannot penalize the distorted or unnatural outputs. To mimic the standard of human visual inspection, we additionally employ Aesthetic score [Schuhmann(2022)], Pickscore [Kirstain et al.(2024)Kirstain, Polyak, Singer, Matiana, Penna, and Levy], and Imagereward [Xu et al.(2024)Xu, Liu, Wu, Tong, Li, Ding, Tang, and Dong] which measure how much people like the image on average.
Figure 3: Qualitative comparison with state-of-the-art methods on Stable Diffusion v1.4.
5.2 Qualitative Evaluation
@@
We compare the editing results of our method to state-of-the-art methods using a pre-trained Stable Diffusion v1.4 in Figure 3. We show our results using Null-text inversion as it provides the best editing performance (Figure 5). For a fair comparison, we also applied Null-text inversion in Masactrl, which originally used DDIM inversion. It results in improved identity preservation and editability of Masactrl. We tried various hyperparameters of each method and chose the best result for each.
@@
Firstly, Null-text inversion alone for editing, in which we just input an inverted latent and a target text through the model, shows low identity preservation. It does not retain the original structure of the input image and simply pursues generating an image that aligns well with the text prompt. This is expected because there is no other constraint for identity preservation during the sampling process. Meanwhile, Imagic has a serious tradeoff between overfitting and editability. Model finetuning especially causes a severe overfitting problem, which distorts the data distribution, whereas a random noise without finetuning during the sampling process ruins the structure preservation. Despite the extensive hyperparameter search, it is hard to find a sweet spot between these two factors (See Appendix). Besides, DDS also suffers from overfitting to the input image, especially when handling non-rigid editing. This shows that score distillation has trouble changing geometry. On the other hand, Edit-Friendly and Masactrl attempt to apply an editing concept while maintaining the original identity by improving latent noise space and utilizing attention maps, respectively. When an image has no complicated background and contains a single object, such as a Spiderman image in the third row of Figure 3, their editing quality is sufficiently high. However, they show unnatural edited results in other cases. In the first row of Figure 3, a bulldog opens its mouth but does not smile in Edit-Friendly, and the bulldog shows a grotesque mouth in Masactrl. In the second row, a horse jumps in both methods, but the hind legs of the horse remain similar to the input image, resulting in an unnatural jumping appearance. Overall, our method clearly shows the editing concept while faithfully preserving the original identity, like a bulldog with a big smile and a horse jumping dynamically. These visual trends and the superiority of our method consistently hold in other results. More qualitative results are demonstrated in Appendix, including results using the pretrained Anything-v4 model.
@@ 5.3 Quantitative Evaluation
Due to inherent inaccuracies and ambiguities in quantitative evaluation of non-rigid editing, we conduct a user study. We utilized the TedBench dataset [Kawar et al.(2023)Kawar, Zada, Lang, Tov, Chang, Dekel, Mosseri, and Irani] to generate 50 edited images of each baseline model. A set of input real image, target text and 6 edited images using each model is shown to the user. The user is asked to choose one among edited images that best satisfies both two conditions simultaneously: The edited image 1) maintains background and appearance of the object in the original input image and 2) aligns with the target text. We randomly shuffled the order of edited images so that the user doesn’t know a model used to generate each image. We accumulated 38 respondents, resulting in total 1900 (38 $\times$ 50) responses. Figure 4 (A) shows the preference rate of each method, which represents the proportion of times the corresponding method was selected out of 1900 responses. Ours receives the highest preference rate, 0.43, which is more than twice as high as the preference rate of the second-ranked Masactrl (0.19). It indicates that our method surpasses other models in terms of both similarity to the input image and text alignment.
@@
We further conduct a quantitative evaluation on the models that ranked within the top three in user study: Edit-Friendly, Masactrl, and Ours. We utilized the same target text prompts used in user study to generate 50 edited images, producing five variants for each edited image with different seeds, resulting in a total of 250 images. In Figure 4 (B), we present various metrics of each method at different CLIP-T scores, where CLIP-T changes depending on hyperparameters. Generally, Edit-Friendly receives the best score in LPIPS ($\downarrow$), PSNR ($\uparrow$) and MSE ($\downarrow$) showing that it has a strong tendency to preserve the identity of the original image. Compared to Edit-Friendly, Ours and Masactrl edit images in a way that align more closely with the text prompt. For non-rigid editing which involves changes in motion of objects, big changes in the structure of objects should be allowed. Therefore, strictly preserving the structure of the original image disturbs editing. Our user study proves that our method well balances input image similarity and editability. This success can be attributed to our inversion and timestep-aware text injection sampling, which preserve the original identity without disrupting the data distribution. This enables text optimization to smoothly transition towards the desired editing concept while maintaining the background and object identity faithfully. Also, our method receives the highest score in Aesthetic score ($\uparrow$) and Imagereward ($\uparrow$) indicating that our edited images are more preferred in terms of human preference. Masactrl, on the other hand, presents the lowest score in Aesthetic score, Pickscore and Imagereward in general, consistent with its tendency to generate unnatural images with distorted object appearances and poses. This can be attributed to the difference in object appearance and poses between the source and target images, each generated using the original source and target text, respectively. Example images are available in Appendix. As the appearance gap between the target image and the source image intensifies, inserting the attention maps of the source image into the target image sampling process often results in abnormal objects with unnatural appearances and poses.
Figure 4: User study and quantitative comparison.
@@ 5.4 Ablation Experiments
Inversion Type. Figure 5 demonstrates the impact of different inversion methods: DDIM [Song et al.(2020)Song, Meng, and Ermon] and Null-text inversion [Mokady et al.(2023)Mokady, Hertz, Aberman, Pritch, and Cohen-Or]. Each image is edited with the target text “A photo of a jumping dog.” When employing DDIM inversion for our method, it is highly close to the input image but has low editability. This again proves that DDIM inversion is suboptimal for real image editing applications. On the other hand, Null-text inversion is known to improve reconstruction and editing quality, particularly when using classifier-free guidance with a guidance scale greater than 1, as compared to DDIM inversion. Interestingly, when inputting a source prompt (e.g. “A photo of a dog") following the original method, the preservation quality sometimes significantly deteriorates while focusing on editing (NTI - src). Conversely, when we set a source text prompt equal to null text (‘’), the edited images show significant improvement in structural preservation (NTI - $\emptyset$). Therefore, when the reconstruction performance falls behind, employing the null text for the source prompt can enhance the overall reconstruction ability. Overall, Null-text inversion demonstrates superior performance in editing quality to DDIM.
Figure 5: Ablation on inversion type.